The sum of two consecutive even numbers is 622. Find the numbers
Answer by HyperBrain(694)

notice that the difference of any two consecutive even (or odd) integers is always 2. evidence----2 and 4 are consecutive even integers and 4-2=2.
So, if x is the smaller number, then x+2 is the larger number
And, their sum is
x+(x+2)
Given that the sum of two consecutive even numbers is 622, then
x+(x+2)=622
(x+x)+2=622
2x+2=622
2x+2-2=622-2
2x+0=620
2x=620
x=310
As expected, x is an even number
So,
x+2=310+2=312

Thus,
if the sum of two consecutive even numbers is 622, then the numbers are 310 and 312.

Power up,
HyperBrain!